OpenStudy (anonymous):

trig proof help pls (sinx / 1 – cosx) + (sinx / 1 + cosx) = 2 csc x

OpenStudy (anonymous):

@ganeshie8 I know you can do these, so can you help?

OpenStudy (anonymous):

\[\frac{ \sin x }{ 1 - \cos x }\frac{ \sin x }{ 1 + \cos x } = 2 \csc x\]

ganeshie8 (ganeshie8):

\(\dfrac{ \sin x }{ 1 - \cos x }+\dfrac{ \sin x }{ 1 + \cos x }\) \(\sin x\left(\dfrac{ 1 }{ 1 - \cos x }+\dfrac{ 1}{ 1 + \cos x }\right)\) see if you can add the fractions inside the parenthesis and simplify a bit..

OpenStudy (anonymous):

sin x (1 - sec x + 1 + sec x) does this work???

OpenStudy (anonymous):

sin x (1 / 2 cos x) maybe??? @ganeshie8

OpenStudy (anonymous):

As for your question, SinX/(1 -- CosX) + (1 -- CosX)/SinX = 2CscX First, we obtain a common denominator. So we multiply the first fraction by SinX, & the 2nd fraction by (1 -- CosX). We get: (SinX*SinX)/[SinX(1 -- CosX)] + (1 -- CosX)(1 -- CosX)/[SinX(1 -- CosX)] = 2CscX [(SinX)^2]/[SinX(1 -- CosX)] + (1 -- CosX)(1 -- CosX)/[SinX(1 -- CosX)] = 2CscX Here we use the Related Angle Identity: (SinX)^2 + (CosX)^2 = 1 [1 -- (CosX)^2]/[SinX(1 -- CosX)] + (1 -- CosX)(1 -- CosX)/[SinX(1 -- CosX)] = 2CscX (1 -- CosX)(1 + CosX)/[SinX(1 -- CosX)] + (1 -- CosX)(1 -- CosX)/[SinX(1 -- CosX)] = 2CscX (1 + CosX)/SinX + (1 -- CosX)(1 -- CosX)/[SinX(1 -- CosX)] = 2CscX Next, we cancel the (1 -- CosX) from the 2nd fraction, as it's obviously un-necessary, we get: (1 + CosX)/SinX + (1 -- CosX)/SinX = 2CscX Since both fractions have the same denominator now, we add the numerators. We get: 2/SinX = 2CscX The CosX terms obviously cancel out. 2CscX = 2CscX Remember that CscX = 1/SinX. Therefore, the equality holds true: SinX/(1 -- CosX) + (1 -- CosX)/SinX = 2CscX I sure hope that was clear & helpful enough. :) Good luck, take care, & have a great day. :)

OpenStudy (anonymous):

one problem...

OpenStudy (anonymous):

what is it huh :(

OpenStudy (anonymous):

for the second fraction you are using (1 - CosX)/SinX when it should be sin x / (1 + cos x)

OpenStudy (anonymous):

does negating allow you to invert it?

OpenStudy (anonymous):

yes that's the only way

OpenStudy (anonymous):

alright, since I can do that, this problem makes a lot more sense. Thank You!!!

OpenStudy (anonymous):

your welcome :)

ganeshie8 (ganeshie8):

Hey we cannot break fractions like that @snackshack79 we need to follow some rules

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

does mjmahmood 's explanation look alright, or is that what you are saying is wrong?

OpenStudy (anonymous):

sorry, one question @ganeshie8 what is going on between steps one and two? it makes sense, but I 'm not sure what you are multiplying by